can u help me with 2(44+-9999)=6n
Answer by edjones(7311)

2(44+-9999)=6n
we have 2 problems let's do one at a time.
2(44+9999)=6n
2*10043=6n
20086=6n
n=20086/6
n= 3,347 2/3
.
2(44-9999)=6n
2*-9955=6n
-19910=6n
n=-19910/6
n= -3,318 1/3
.
Answer n= 3,347 2/3, n= -3,318 1/3
.
Ed